# Cancellation properties for exotic $4$-dimensional positive scalar curvature metrics

### v2

@@ 2. Proof of Theorem A
Suppose that $M_{1}$ is a $\theta$-manifold having a psc metric and such that $[M_{0}]=[M_{1}]\in\Omega_{4}^{\theta}$. By surgery below the middle dimension (which works in dimension $4$ without change), $[M_{1}]$ contains a representative $[M_{2}]$ such that the structure map $M_{2}\to B$ is $2$-connected. Since we can obtain $M_{2}$ from $M_{1}$ by doing surgeries along embedded spheres of dimension $\leq 1$, i.e. of codimension $\geq 3$, the Gromov–Lawson surgery theorem implies that $M_{2}$ carries a psc metric as well. Theorem C of [27] shows that $M_{2}\sharp\ell(S^{2}\times S^{2})\cong M_{0}\sharp k(S^{2}\times S^{2})$ for suitable $k,\ell$. Since $M_{2}$ has a psc metric, so does $M_{2}\sharp\ell(S^{2}\times S^{2})\cong M_{0}\sharp k(S^{2}\times S^{2})$, proving claim (1).
For claim (2), let $M_{2}$ be as above and note that $M_{2}\times N$ carries a psc metric (by shrinking the metric on $M_{2}$). Let $B_{N}$ be the normal $2$-type of $N$. The structure maps $M_{i}\times N\to B\times B_{N}$ are $2$-connected, and when $\theta\times\theta_{N}:B\times B_{N}\to BO\times BO\times BO$ denotes the sum of the two tangential structures, we have $[M_{0}\times N]=[M_{2}\times N]\in\Omega_{4+\dim(N)}^{\theta\times\theta_{N}}$. Now [11, Theorem 1.5] implies that $M_{0}\times N$ has a psc metric. ∎
@@ 3. Rigidity theorem for the diffeomorphism action in high dimensions
In this section, we describe the aforementioned rigidity theorems and show how they are used in the proof of Theorem B. This will lead to a cobordism-theoretic criterion concerning $f$; if $n\geq 2$, this criterion is easily check, the relevant calculation in the case $n=1$ is carried out in §4.
@@ 3.1. Frenck’s theorem
The surprising result from [16] is that when $\dim(M)\geq 6$, even though source and target of (3.1) are complicated/intractable, the homomorphisms (3.1) itself is relatively easy to control in terms of cobordism theory.
The precise statement requires to use tangentially structured diffeomorphisms. Let $B\stackrel{{\scriptstyle\theta}}{{\to}}BO$ be a map, and let us write $\theta(n):B(n)\to BO(n)$ for the homotopy pullback along $BO(n)\to B$ and $V^{\theta}_{n}\to B(n)$ for the pullback of the universal vector bundle along $\theta(n)$.
For a $d$-manifold $M$, we write $\mathrm{Bun}(TM,\theta)$ for the space of bundle maps (fibrewise isomorphisms) $TM\to V^{\theta}_{d}$, with the compact-open topology. We let
$B\mathrm{Diff}^{\theta}(M):=E\mathrm{Diff}(M)\times_{\mathrm{Diff}(M)}\mathrm{Bun}(TM,\theta).$
This classifies smooth $M$-bundles $\pi:E\to X$, together with a bundle map $T_{v}E\to V_{d}^{\theta}$ from its vertical tangent bundle. We mention that $B\mathrm{Diff}^{\theta}(M)$ is usually disconnected and hence not the classifying space of a topological group; in [3, Examples 1.32] we worked out the homotopy type in some simple cases.
As explained in [3, §1.7], it is conceptually useful to think about $B\mathrm{Diff}^{\theta}(M)$ as the space of all $(N,\ell)$, where $N$ is a closed manifold which is diffeomorphic to $M$ and $\ell:TN\to V_{d}^{\theta}$ is a $\theta$-structure on $N$. In light of this, a point $\ell\in\mathrm{Bun}(TM,\theta)$ gives rise to a basepoint $(M,\ell)\in B\mathrm{Diff}^{\theta}(M)$, and we define
@@
If the precise choice of $\ell$ does not matter, we just write $\Gamma^{\theta}(M)$. Forgetting the bundle data gives map $B\mathrm{Diff}^{\theta}(M)\to B\mathrm{Diff}(M)$ inducing on fundamental groups a homomorphism
@@
Two cases deserve special attention.
@@
    If $\theta:B\mathrm{Spin}\to BO$, $\ell$ corresponds to a spin structure on $M$. Then (3.2) has kernel of order $2$ and its image consists of those mapping classes which fix $\ell$ up to isomorphism. If in addition $M$ is $1$-connected, the image of (3.2) agrees with $\Gamma^{+}(M)$.
Now assume that $f\in\mathrm{Diff}(M)$ is such that $[f]\in\Gamma(M)$ lies in the image of (3.2). A choice of a preimage in $\Gamma^{\theta}(M,\ell)$ gives rise to a bundle map $T_{v}T(f)\to V_{d}^{\theta}$ from the vertical tangent bundle of the mapping torus. Since $TT(f)=T_{v}T(f)\oplus\mathbb{R}$, we get a bundle map $TT(f)\to V_{d+1}^{\theta}$, which gives the manifold $T(f)$ a $\theta$-structure. We therefore get an element $[T(f)]\in\Omega_{d+1}^{\theta}$, the bordism group of $(d+1)$-manifolds with stable tangential $\theta$-structure [^2] Examples (for $B=B\mathrm{Spin}$) show that $[T(f)]$ actually depends on the choice of the lift of $[f]$ along (3.2). This won’t cause any trouble for us.. This construction gives a map
@@
generalizing (1.7). As (1.7), (3.4) is a group homomorphism (see [16, Corollary 3.30] for the proof in the general case). With these preliminaries in place, we quote from [16]:
Theorem 3.5 (Frenck).
Let $M$ be a closed manifold of dimension $d\geq 6$ and let $\ell\in\mathrm{Bun}(TM,\theta)$ be so that the underlying map $M\to B$ is $2$-connected. Then there exists a group homomorphism $F:\Omega^{\theta}_{d+1}\to\pi_{0}(\mathrm{hAut}(\mathcal{R}^{+}(M)))$ such that the diagram
@@
The main ingredient for Theorem 3.5 is the Gromov–Lawson–Chernysh surgery theorem [11, Theorem 1.2] and Igusa’s $2$-index theorem [26, Chapter VI].
Theorem 3.5 was preceeded by [4, Theorem 4.1.2] and [12, Theorem 4.1.2]. Both these results established (for some special $M$) diagrams similar to the one in 3.5, but with an a priori unknown abelian group in place of $\Omega^{\theta}_{d+1}$.
@@ 3.2. Bantje’s theorem
The case $n=1$ of Theorem B requires us to use the more involved (to prove, state and use) rigidity result [2, Theorem 10.4.1]. The latter is stronger that Theorem 3.5 in three aspects:
@@
    it applies to manifolds with nonempty boundary.
However, in the case where both results apply, Theorem 3.5 has a slightly stronger conclusion (reflected in the stronger hypotheses of Theorem B in the case $n=1$). We will not state the version of [2, Theorem 10.4.1] for manifolds with boundary here. Having given this brief overview, let us turn to the actual statement, quoting [2, Theorem 10.4.1] (and explaining its constitutents afterwards).
@@
Let $M$ be a closed manifold of dimension $d\geq 5$ and let $\theta:B\to BO$ be a fibration. Then there is a homotopy cartesian diagram
@@
    The left column of (3.7) is just the Borel construction $\mathcal{R}^{+}(M)\sslash\mathrm{Diff}(M):=E\mathrm{Diff}(M)\times_{\mathrm{Diff}(M)}\mathcal{R}^{+}(M)\to B\mathrm{Diff}(M)$ coming from the diffeomorphism action.
@@
    $E\mathrm{Diff}(M)\times_{\mathrm{Diff}(M)}\mathrm{Bun}(TM,\theta)^{(2)}$
    The left horizontal map of (3.7) just forgets the bundle data, and the left square of (3.7) is defined to be a pullback. In other words
@@
    The space $\mathfrak{X}$ is constructed in [2] and (of course) the main point of the theorem. The upper right horizontal map is where the condition of the $2$-connectivity comes in.
@@
    If either $\mathcal{R}^{+}(M)=\emptyset$ or $\mathrm{Bun}(TM,\theta)^{(2)}=\emptyset$, the result is vacuously true (hence a sensible choice of $\theta$ is crucial when applying the result).
@@
    The unstated version of Theorem 3.6 for manifolds with boundary supersedes [13, Theorems E and F] which apply only for $d\geq 6$ and exclusively to manifolds with nonempty boundary. The latter result implies the main result of [5] which might be viewed as the first instance of result in this direction. With hindsight, these theorems have been the main obstacle in the attempt to directly use the topology of $\mathrm{Diff}(M)$ to get information about $\mathcal{R}^{+}(M)$ [25], [7], [22].
@@
    In some special case, diagrams similar to (3.7) have been established in [4, Diagram (1.6)] and [12, Theorem F], as consequences of the special cases of Theorem 3.5 proved in those papers. These have been the decisive steps in those papers.
  3. (3)
    Let us mention that the proof of Theorem 3.6 relies on ideas of Perlmutter [34], [35] which in turn are an elaboration of parts of the original proof of the Mumford conjecture by Madsen and Weiss [29] to high dimensions. Of course, the Gromov–Lawson–Chernysh theorem is an irreplacable ingredient of Theorem 3.6.
Let us next explain how to deduce from Theorem 3.6 a statement that is formally similar to that of Theorem 3.5. The left column of (3.7) is by definition a fibration with fibre $\mathcal{R}^{+}(M)$. As such, it has a classifying map $B\mathrm{Diff}(M)\to B\mathrm{hAut}(\mathcal{R}^{+}(M))$, by the general classification theory for fibrations [30]. Of course, this map just comes from delooping the action map $\mathrm{Diff}(M)\to\mathrm{hAut}(\mathcal{R}^{+}(M))$ and does not carry any new information. The composition $B\mathrm{Diff}^{\theta}(M)^{(2)}\to B\mathrm{Diff}(M)\to B\mathrm{hAut}(\mathcal{R}^{+}(M))$ is then, tautologically, a classifying map for the middle fibration in (3.7). Now comes the point: since the right square of (3.7) is homotopy cartesian, the classifying map for the middle map can be factored through $\Omega^{\infty}\mathrm{MT}\theta(d)$.
The uniqueness up to homotopy of these classifying maps therefore gives from (3.7) a homotopy commutative diagram
${B\mathrm{Diff}^{\theta}(M)^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${\Omega^{\infty}\mathrm{MT}\theta(d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${B\mathrm{Diff}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\mathrm{hAut}(\mathcal{R}^{+}(M)).}$ (3.10)
Now let $\ell:TM\to V_{d}^{\theta}$ be a bundle map whose underlying map $M\to B(d)$ is $2$-connected. Taking fundamental groups at $\ell\in B\mathrm{Diff}^{\theta}(M)^{(2)}$ in (3.10) gives a commutative diagram [^3] The Madsen–Tillmann map $\alpha:B\mathrm{Diff}^{\theta}(M)\to\Omega^{\infty}\mathrm{MT}\theta(d)$ sends $(M,\ell)$ to a $\alpha(M,\ell)\in\Omega^{\infty}\mathrm{MT}\theta(d)$, which is the basepoint for the fundamental group in the upper right corner.
@@
of groups. Since $\Omega^{\infty}\mathrm{MT}\theta(d)$ is an infinite loop space, the fundamental groups at $\alpha(M)$ is canonically identified with the fundamental group at the basepoint which in turn is identified with the homotopy group $\pi_{1}(\mathrm{MT}\theta(d))$ of the spectrum $\mathrm{MT}\theta(d)$ (keeping track of this identification will require some care further below).

### v3

@@ 2. Proof of Theorem A
Suppose that $M_{1}$ is a $\theta$-manifold having a psc metric and such that $M_{0}]=[M_{1}]\in\Omega_{4}^{\theta}$. By surgery below the middle dimension (which works in dimension $4$ without change), $M_{1}]$ contains a representative $M_{2}]$ such that the structure map $M_{2}\to B$ is $2$-connected. Since we can obtain $M_{2}$ from $M_{1}$ by doing surgeries along embedded spheres of dimension $\leq 1$, i.e. of codimension $\geq 3$, the Gromov–Lawson surgery theorem implies that $M_{2}$ carries a psc metric as well. Theorem C of [30] shows that $M_{2}\sharp\ell(S^{2}\times S^{2})\cong M_{0}\sharp k(S^{2}\times S^{2})$ for suitable $k,\ell$. Since $M_{2}$ has a psc metric, so does $M_{2}\sharp\ell(S^{2}\times S^{2})\cong M_{0}\sharp k(S^{2}\times S^{2})$, proving claim (1).
For claim (2), let $M_{2}$ be as above and note that $M_{2}\times N$ carries a psc metric (by shrinking the metric on $M_{2}$). Let $B_{N}$ be the normal $2$-type of $N$. The structure maps $M_{i}\times N\to B\times B_{N}$ are $2$-connected, and when $\theta\times\theta_{N}:B\times B_{N}\to BO\times BO$ denotes the sum of the two tangential structures, we have $M_{0}\times N]=[M_{2}\times N]\in\Omega_{4+\dim(N)}^{\theta\times\theta_{N}}$. Now [14, Theorem 1.5] implies that $M_{0}\times N$ has a psc metric. ∎
@@ 3. Rigidity theorem for the diffeomorphism action in high dimensions
In this section, we describe the aforementioned rigidity theorems and show how they are used in the proof of Theorem B. This will lead to a cobordism-theoretic criterion concerning $f$; if $n\geq 2$, this criterion is easily checked, and the relevant calculation in the case $n=1$ is carried out in §4.
@@ 3.1. Frenck’s theorem
The surprising result from [19] is that when $\dim(M)\geq 6$, even though source and target of (3.1) are complicated/intractable, the homomorphism (3.1) itself is relatively easy to control in terms of cobordism theory.
The precise statement involves the notion tangentially structured diffeomorphisms which we briefly recall. Let $\theta:B\to BO$ be a map, and let us write $\theta(n):B(n)\to BO(n)$ for the homotopy pullback along $BO(n)\to BO$ and $V^{\theta}_{n}\to B(n)$ for the pullback of the universal vector bundle along $\theta(n)$.
Definition 3.2.
A $\theta$-structure on a $d$-manifold $M$ is a bundle map (fibrewise isomorphisms) $\ell:TM\to V^{\theta}_{d}$. We write $\mathrm{Bun}(TM,\theta)$ for the space of bundle maps $TM\to V^{\theta}_{d}$, with the compact-open topology and define
$B\mathrm{Diff}^{\theta}(M):=E\mathrm{Diff}(M)\times_{\mathrm{Diff}(M)}\mathrm{Bun}(TM,\theta)\to B\mathrm{Diff}(M).$ (3.3)
This classifies smooth $M$-bundles $\pi:E\to X$, together with a bundle map $T_{v}E\to V_{d}^{\theta}$ from its vertical tangent bundle; the map (3.3) forgets the bundle data. We mention that $B\mathrm{Diff}^{\theta}(M)$ is often empty or disconnected and hence not the classifying space of a topological group; in [6, Examples 1.32] we worked out the homotopy type in some simple cases.
As explained in [6, §1.7], it is conceptually useful to think about $B\mathrm{Diff}^{\theta}(M)$ as the space of all $(N,\ell)$, where $N$ is a closed manifold which is diffeomorphic to $M$ and $\ell:TN\to V_{d}^{\theta}$ is a $\theta$-structure on $N$. In light of this, a point $\ell\in\mathrm{Bun}(TM,\theta)$ gives rise to a basepoint $(M,\ell)\in B\mathrm{Diff}^{\theta}(M)$, and we define
@@
If the precise choice of $\ell$ does not matter, we just write $\Gamma^{\theta}(M)$. The map (3.3) induces on fundamental groups a homomorphism
@@
Three cases deserve special attention.
  1. (1)
    If $\theta:BO\to BO$ is the identity, $\mathrm{Bun}(TM,\theta)\simeq*$, so (3.4) is an isomorphism.
@@
    If $\theta:B\mathrm{Spin}\to BO$, $\ell$ corresponds to a spin structure on $M$. Then (3.4) has kernel of order 22 and its image consists of those mapping classes which fix $\ell$ up to isomorphism. If in addition $M$ is $1$-connected, the image of (3.4) agrees with $\Gamma^{+}(M)$.
Now assume that $f\in\mathrm{Diff}(M)$ is such that $f]\in\Gamma(M)$ lies in the image of (3.4). A choice of a preimage in $\Gamma^{\theta}(M,\ell)$ gives rise to a bundle map $T_{v}T(f)\to V_{d}^{\theta}$ from the vertical tangent bundle of the mapping torus. Since $TT(f)=T_{v}T(f)\oplus\mathbb{R}$, we get a bundle map $TT(f)\to V_{d+1}^{\theta}$, which gives the manifold $T(f)$ a $\theta$-structure. We therefore get an element $T(f)]\in\Omega_{d+1}^{\theta}$, the bordism group of $(d+1)$-manifolds with stable tangential $\theta$-structure [^2] Examples (for $B=B\mathrm{Spin}$) show that $[T(f)]$ actually depends on the choice of the lift of $[f]$ along (3.4). This won’t cause any trouble for us.. This construction gives a map
@@
generalizing (1.7). Just as the map (1.7), the more general map (3.6) is a group homomorphism (see [19, Corollary 3.30] for the proof in the general case).
Theorem 3.7 (Frenck [19]).
Let $M$ be a closed manifold of dimension $d\geq 6$ and let $\ell\in\mathrm{Bun}(TM,\theta)$ be a $2$-connected $\theta$-structure on MM [^3] In other words, the map $M\to B$ underlying $\ell$ is $2$-connected. Then there exists a group homomorphism $F:\Omega^{\theta}_{d+1}\to\pi_{0}(\mathrm{hAut}(\mathcal{R}^{+}(M)))$ such that the diagram
@@
The main ingredients for Theorem 3.7 are the Gromov–Lawson–Chernysh surgery theorem [14, Theorem 1.2] and Igusa’s $2$-index theorem [29, Chapter VI].
Theorem 3.7 was preceeded by [7, Theorem 4.1.2] and [15, Theorem 4.1.2]. Both these results established (for some special $M$) diagrams similar to the one in 3.7, but with an a priori unknown abelian group in place of $\Omega^{\theta}_{d+1}$.
Remark 3.8.
Let $f]\in\Gamma^{\theta}(M,\ell)$. We emphasize that the theorem gives a sufficient condition for $f^{*}\sim\mathrm{id}\in\mathrm{hAut}(\mathcal{R}^{+}(M))$: if the map underlying $\ell$ is $2$-connected and if $[f]$ maps to $0\in\Omega_{d+1}^{\theta}$ under (3.6), then $f^{*}\simeq\mathrm{id}\in\mathrm{hAut}(\mathcal{R}^{+}(M))$.
A necessary condition for $f^{*}\simeq\mathrm{id}$ is as follows. Regardless of $d\geq 6$ and $\ell$ being $2$-connected or not, $f^{*}\sim\mathrm{id}$ implies that the mapping torus $T(f)$ carries a psc metric. It follows that the image of $[f]$ under (3.6) lies in the subgroup $(\Omega_{d+1}^{\theta})_{+}\subset\Omega_{d+1}^{\theta}$ of cobordism classes having a psc representative.
Therefore for $f]\in\Gamma^{\theta}(M,\ell)$, we deduce
$T(f)]=0\in\Omega_{d+1}^{\theta}\stackrel{{\scriptstyle d\geq 6,\ell\,2-\text{connected}}}{{\Rightarrow}}f^{*}\sim\mathrm{id}\Rightarrow[T(f)]\in(\Omega_{d+1}^{\theta})_{+}.$
In the case of $\theta:B\mathrm{Spin}\to BO$, the group $(\Omega_{d+1}^{\mathrm{Spin}})_{+}\subset\Omega_{d+1}^{\mathrm{Spin}}$ coincides with the kernel of the $\alpha$-invariant $\hat{\mathfrak{a}}:\Omega_{d+1}^{\mathrm{Spin}}\to\mathrm{KO}_{d+1}$, by the main result of [44]. Usually $(\Omega_{d+1}^{\mathrm{Spin}})_{+}$ is not the trivial subgroup.
More generally, one can phrase the necessary condition more succinctly in terms of the Stolz exact sequence from [43]
$\ldots\to\Omega_{d+1}^{\theta,\mathrm{psc}}\stackrel{{\scriptstyle F}}{{\to}}\Omega_{d+1}^{\theta}\stackrel{{\scriptstyle G}}{{\to}}R_{d+1}^{\theta}\stackrel{{\scriptstyle\partial}}{{\to}}\Omega_{d}^{\theta,\mathrm{psc}}\to\ldots.$
Here $\Omega_{d}^{\theta,\mathrm{psc}}$ is the cobordism group of closed $d$-dimensional $\theta$-manifolds with psc metrics (and psc cobordisms); $F$ forgets the psc metrics and $(\Omega_{d+1}^{\theta})_{+}=\mathrm{im}(F)$; $R_{d+1}^{\theta}$ is a relative group of compact $(d+1)$-dimensional $\theta$-manifolds with psc metrics on the boundary and $\partial$ takes the boundary. So for $f]\in\Gamma^{\theta}(M,\ell)$, we deduce
$T(f)]=0\in\Omega_{d+1}^{\theta}\stackrel{{\scriptstyle d\geq 6,\ell\,2-\text{connected}}}{{\Rightarrow}}f^{*}\sim\mathrm{id}\Rightarrow G([T(f)])=0\in R_{d+1}^{\theta}.$
One might speculate whether one of the above implications is an equivalence, but that is a wide open question.
@@ 3.2. Bantje’s theorem
The case $n=1$ of Theorem B requires us to use the more involved (to prove, state and use) rigidity result [5, Theorem 10.4.1], which is stronger that Theorem 3.7 in three aspects.
@@
    it applies to manifolds with nonempty boundary (we won’t discuss this here).
However, in the case where both results apply, Theorem 3.7 has a slightly stronger conclusion (reflected in the stronger hypotheses of Theorem B in the case $n=1$). Having given this brief overview, let us turn to the actual statement, quoting [5, Theorem 10.4.1] (and explaining its constitutents afterwards).
@@
Let $M$ be a closed manifold of dimension $d\geq 5$ and let $\theta:B\to BO$ be a map. Then there is a homotopy cartesian diagram
@@
    The left column of (3.10) is just the Borel construction
    $\mathcal{R}^{+}(M)\sslash\mathrm{Diff}(M):=E\mathrm{Diff}(M)\times_{\mathrm{Diff}(M)}\mathcal{R}^{+}(M)\to B\mathrm{Diff}(M)$
    coming from the diffeomorphism action.
@@
    $B\mathrm{Diff}^{\theta}(M)^{(2)}:=E\mathrm{Diff}(M)\times_{\mathrm{Diff}(M)}\mathrm{Bun}(TM,\theta)^{(2)}.$
    The lower left horizontal map of (3.10) is the restriction of (3.3), and the left square of (3.10) is defined to be a pullback. In other words
@@
    The space $\mathfrak{X}$ is constructed in [5] and (of course) the main point of the theorem, and also the place where the 22-connectivity condition enters. In §3.3 below, we give a brief sketch.
@@
    If either $\mathcal{R}^{+}(M)=\emptyset$ or $\mathrm{Bun}(TM,\theta)^{(2)}=\emptyset$, the result is vacuously true (with $\mathfrak{X}=\emptyset$). Hence a sensible choice of $\theta$ is crucial when applying the result.
@@
    The unstated version of Theorem 3.9 for manifolds with boundary supersedes [16, Theorems E and F] which apply only for $d\geq 6$ and exclusively to manifolds with nonempty boundary. The latter result implies the main result of [8] which might be viewed as the first instance of result in this direction. With hindsight, these theorems have been the main obstacle in the attempt to directly use the topology of $\mathrm{Diff}(M)$ to get information about $\mathcal{R}^{+}(M)$ [28], [10], [25].
@@
    In some special cases, diagrams similar to (3.10) have been established in [7, Diagram (1.6)] and [15, Theorem F], as consequences of the special cases of Theorem 3.7 proved in those papers. These have been the decisive steps in those papers.
Let us next explain how to deduce from Theorem 3.9 a statement that is formally similar to that of Theorem 3.7. A $2$-connected $\theta$-structure $\ell\in\mathrm{Bun}(TM,\theta)^{(2)}$ on $M$ determines $(M,\ell)\in B\mathrm{Diff}^{\theta}(M)^{(2)}$; let $B\mathrm{Diff}^{\theta}(M)^{(2)}_{(M,\ell)}$ be the path component containing that basepoint. We restrict (3.10) to obtain the diagram
${\mathcal{R}^{+}(M)\sslash\mathrm{Diff}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\mathrm{Diff}^{\theta,\mathrm{psc}}(M)^{(2)}_{(M,\ell)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathfrak{X}_{(M,\ell)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\mathrm{Diff}(M)}$${B\mathrm{Diff}^{\theta}(M)^{(2)}_{(M,\ell)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${\Omega^{\infty}_{(M,\ell)}\mathrm{MT}\theta(d);}$ (3.13)
here $\Omega^{\infty}_{(M,\ell)}\mathrm{MT}\theta(d)\subset\Omega^{\infty}\mathrm{MT}\theta(d)$ is the path component containing the image of $(M,\ell)$ under $\alpha$. The upper row is obtained by taking the preimages of the distinguished path components; hence the diagram (3.13) is still homotopy cartesian.
The left column of (3.13) is by definition a fibration with fibre $\mathcal{R}^{+}(M)$. As such, it has a classifying map $B\mathrm{Diff}(M)\to B\mathrm{hAut}(\mathcal{R}^{+}(M))$, unique up to homotopy, by the general classification theory for fibrations [33]. Of course, this map just comes from delooping the action map $\mathrm{Diff}(M)\to\mathrm{hAut}(\mathcal{R}^{+}(M))$ and does not carry any new information. The composition $B\mathrm{Diff}^{\theta}(M)^{(2)}_{(M,\ell)}\to B\mathrm{Diff}(M)\to B\mathrm{hAut}(\mathcal{R}^{+}(M))$ is then, tautologically, a classifying map for the middle fibration in (3.13). The fact that the right square in (3.13) is homotopy cartesian implies that the classfying map for the middle fibration factors, up to homotopy, through $\Omega^{\infty}_{(M,\ell)}\mathrm{MT}\theta(d)$. Therefore we establish from (3.13) a homotopy commutative diagram
${B\mathrm{Diff}^{\theta}(M)^{(2)}_{(M,\ell)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha}$${\Omega^{\infty}_{(M,\ell)}\mathrm{MT}\theta(d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi}$${B\mathrm{Diff}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\mathrm{hAut}(\mathcal{R}^{+}(M)).}$ (3.14)
Taking fundamental groups at $(M,\ell)\in B\mathrm{Diff}^{\theta}(M)^{(2)}$ and its image points under the maps in (3.14) gives a commutative diagram [^4] The Madsen–Tillmann map $\alpha:B\mathrm{Diff}^{\theta}(M)\to\Omega^{\infty}\mathrm{MT}\theta(d)$ sends $(M,\ell)$ to a $\alpha(M,\ell)\in\Omega^{\infty}\mathrm{MT}\theta(d)$, which is the basepoint for the fundamental group in the upper right corner.
@@
of groups. Since $\Omega^{\infty}\mathrm{MT}\theta(d)$ is an infinite loop space, the fundamental group at $\alpha(M,\ell)$ is canonically identified with the fundamental group at the basepoint which in turn is identified with the homotopy group $\pi_{1}(\mathrm{MT}\theta(d))$ of the spectrum $\mathrm{MT}\theta(d)$ (keeping track of this identification will require some care further below).
3.3. A very brief overview of the proof of Bantje’s theorem
Let us mention that the proof of Theorem 3.9 relies on ideas of Perlmutter [37], [38] which in turn are an elaboration of large parts of the original proof of the Mumford conjecture by Madsen and Weiss [32] in high dimensions. Given the well-known complexity of [32], the reader should not expect an easy ride.
One first introduces a certain spaces of manifolds $\mathcal{W}_{\theta,T}^{[3]}$ and $\mathcal{W}_{\theta,T}^{[3],\mathrm{psc}}$. Here $T$ is a finite set, equipped with some extra data. The points in $\mathcal{W}_{\theta,T}^{[3]}$ are $d$-dimensional closed manifolds with $\theta$-structures and a collection of surgery data of index in $\{3,\ldots,d-2\}$, indexed by the elements of $T$. The $\theta$-structures are required to be 2=312=3-1-connected. The points of $\mathcal{W}_{\theta,T}^{[3],\mathrm{psc}}$ are the points of $\mathcal{W}_{\theta,T}^{[3]}$, but with psc metrics as additional datum; the psc metrics need to be of standard form on the surgery data.
Almost by definition, the forgetful map $\mathcal{W}_{\theta,\emptyset}^{[3],\mathrm{psc}}\to\mathcal{W}_{\theta,\emptyset}^{[3]}$ is the disjoint union of the middle maps in (3.10), over all closed $M$.
As $T$ ranges through some indexing category $\mathcal{K}^{3}$ of finite sets with extra data, the space $\mathcal{W}_{\theta,T}^{[3]}$ and its psc variant depend functorially on $T$ (by either performing surgeries along the surgery data or by dropping the surgery data), and we can form the homotopy colimit over $\mathcal{K}^{3}$. Together with the natural inclusion maps to the homotopy colimit, one obtains a commutative diagram
${B\mathrm{Diff}^{\theta,\mathrm{psc}}(M)^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{W}_{\theta,\emptyset}^{[3],\mathrm{psc}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{hocolim}_{T\in\mathcal{K}^{3}}\mathcal{W}_{\theta,T}^{[3],\mathrm{psc}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B\mathrm{Diff}^{\theta}(M)^{(2)}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathcal{W}_{\theta,\emptyset}^{[3]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{hocolim}_{T\in\mathcal{K}^{3}}\mathcal{W}_{\theta,T}^{[3]}.}$ (3.16)
The left square is homotopy cartesian almost by definition. The right square is homotopy cartesian as well: this uses a well-known general criterion for proving cartesianness of such squares; the hypothesis of that criterion holds in the case at hand by the Gromov–Lawson–Chernysh theorem.
The crucial point in the proof is to show that the bottom composition in (3.16) can be factored as
$B\mathrm{Diff}^{\theta}(M)^{(2)}\to\Omega^{\infty}\mathrm{MT}\theta(d)\to\mathrm{hocolim}_{T\in\mathcal{K}^{3}}\mathcal{W}_{\theta,T}^{[3]}.$ (3.17)
Once this is done, we define $\mathfrak{X}$ as the homotopy pullback of the right column of (3.16) along the second map of (3.17), and Theorem 3.9 follows. The construction of the factorization (3.17) involves an elaborate parametrized surgery argument, generalizing the more mysterious portions of [32] to high dimensions, and we won’t attempt an informative synopsis here.